OpenStudy (b77w):

Simplify this: (-3x^-2)^-3

OpenStudy (anonymous):

I believe that the answer is -27x^6

OpenStudy (b77w):

How?

OpenStudy (zpupster):

(x)^-n = 1/x^n so in this ex. (-3x^-2)^-3 = (1/-3x^2)^-3 = 1/1/(-3x^2)^3 = 1/1/-27x^6 inverting the fraction and mult. -27x^6

OpenStudy (skullpatrol):

$$(-3x^{-2})^{-3} = (\frac{1}{-3x^2})^{-3} = \frac{1}{\frac{1}{(-3x^2)^3}} = \frac{1}{\frac{1}{-27x^6}} =-27x^6$$

OpenStudy (b77w):

Thanks!